# Posterior SBC: Simulation-Based Calibration Checking Conditional on Data

### v1

@@ 1 Introduction
Figure 1: Illustration of a joint space of model parameters plausible under the priors. Regions 1 and 3 exhibit bias in opposite directions, while inference is well calibrated for parameter values within region 2. Prior SBC will not suggest calibration issues, while posterior SBC can be used to assess trustworthiness for data with posterior contained in one of the regions.
@@
To ensure trust on the results of the inference, different forms of calibration checking play important roles in the Bayesian workflow for model building. Simulation-based calibration checking (SBC) is one of these methods and validates the chosen posterior inference algorithm as well as checks for inconsistencies between the model implementation and a possibly separate implementation of the data generating process. As the software and algorithms for probabilistic modelling advance, the threshold for specifying increasingly complicated models gets lowered. However, as the complexity of the models increases, the probability of human error during the implementation of the model increases. In addition to such implementation mistakes, the inference algorithms used for obtaining posterior approximations can work well for one posterior, but suffer from computational difficulties for another. By simply changing the observed data, one may obtain vastly different posterior geometries, favoring one implementation or algorithm over another.
SBC validates both the implementation of the model and the inference algorithm, but it is limited to checking for calibration on the whole joint distribution of parameters and observations under the specified priors. Specifying prior distributions that accurately present the available prior information is often difficult. In some cases, the prior specifications result in SBC indicating issues, when in truth these issues are limited to a region of the prior space which is of no interest to the modeller. SBC also tends to be computationally expensive, as multiple model refits are required. One can choose to save computational resources by reducing the number of refits, but this in turn also reduces the power of the calibration checking procedure. When reducing the computational budget, SBC may fail to detect problems in the posterior inference if they are infrequent on data drawn from the full prior predictive distribution. After collecting data and running inference, these issues may be present in a large proportion of plausible parameter values.
@@
In this paper, we introduce posterior SBC, a variant of SBC validating the inference algorithm and model implementation when conditioned on some fixed set of data. First, we introduce the method of simulation-based calibration as it has traditionally been operationalized. To emphasize the distinction between these methods, we denote this traditional implementation as prior SBC. Next, in Section 2, we introduce posterior SBC. In Section 3 we illustrate the differences between prior and posterior SBC through three case studies including a simple hierarchical model, Lotka-Volterra model, and a drift-diffusion model. Finally, in section 4 we summarize the contributions of the paper and provide an outlook for future research.
1.1 Simulation-based calibration checking
Cook et al., (2006) proposed a simulation-based calibration checking method for validating Bayesian inference software. The method is based on factoring the joint distribution of data $y$ and model parameters $\theta$ in two ways;
$\pi(y,\theta)=\pi(y\,|\,\theta)\pi(\theta)=\pi(\theta\,|\,y)\pi(y).$ (1)
When considering two draws, $\theta^{\prime},\theta^{\prime\prime}\sim\pi(\theta)$, and data $y$, the joint distribution is
$\pi(y,\theta^{\prime},\theta^{\prime\prime})=\pi(y)\pi(\theta^{\prime}\,|\,y)\pi(\theta^{\prime\prime}\,|\,y).$ (2)
Above, it’s easy to see that $\theta^{\prime}$ and $\theta^{\prime\prime}$ have the same distribution conditionally on $y$. If we use (1) to rewrite the joint distribution in (2) as
$\pi(y,\theta^{\prime},\theta^{\prime\prime})=\pi(\theta^{\prime})\pi(y\,|\,\theta^{\prime})\pi(\theta^{\prime\prime}\,|\,y),$
$\theta^{\prime}$ and $\theta^{\prime\prime}$ still have the same distribution conditionally on $y$.
We can sample from the joint distribution $\pi(y,\theta^{\prime},\theta^{\prime\prime})$ by first sampling $\theta^{\prime}\sim\pi(\theta)$ and $y\sim\pi(y\,|\,\theta^{\prime})$, which is usually easy for generative models. The last step is to sample from the conditional $\pi(\theta^{\prime\prime}\,|\,y)$, which is usually not trivial and requires the use of some inference algorithm, such as Markov chain Monte Carlo (MCMC). We can validate the algorithm and its implementation for sampling $\theta^{\prime\prime}\sim\pi(\theta\,|\,y)$ by checking that, conditionally on $y$, the samples obtained have the same distribution as $\theta^{\prime}$.
@@ 1.1 Simulation-based calibration checking
To assess whether $\theta^{\prime}$ and $\theta^{\prime\prime}$ are from the same distribution conditional on $y$, Cook et al., (2006) repeatedly draw parameter vectors $\theta^{\prime}_{i}$ from the prior $\pi(\theta)$, generate data from the observation model $y_{i}\sim\pi(y_{i}\,|\,\theta^{\prime}_{i})$, and then use the algorithm to be validated to sample from the posterior $\theta^{\prime\prime}_{i,1},\ldots,\theta^{\prime\prime}_{i,S}\sim\pi(\theta^{\prime\prime}\,|\,y_{i})$. They then propose to compute empirical probability integral transform (PIT) values, $u_{i}=p(\theta_{i}^{\prime\prime}<\theta_{i}^{\prime}\,|\,y_{i})$, which prove to be uniformly distributed as $S\to\infty$. The process is repeated for $i=1,\ldots,N$ and the empirical PIT values are used for testing. Figure 2 shows three iterations of prior draws and their respective PIT relative to the corresponding posterior distributions. Cook et al., (2006) further propose the use of the $\chi^{2}$-test for the inverse of the normal cumulative distribution function (CDF) of the empirical PIT values.
@@
Instead of only assessing the uniformity of the ranks of individual parameters, Modrák et al., (2023) propose the use of the joint log-likelihood of the prior predictive sample as a test statistic for assessing the calibration. This has the additional power of detecting calibration issues rising from the model partially ignoring the data. Another advantage of this approach lies in modelling applications with high-dimensional parameter spaces, where modelers might not be particularly interested in gauging the calibration of individual parameters. Here, using the joint log-likelihood avoids the need of correcting for multiple comparisons from simultaneously assessing the calibration of all individual parameters.
@@
Performing SBC on the full prior parameter space has three fundamental drawbacks. First, as shown in the case study of Section 3.1, a region of calibration issues may have relatively small effect on prior SBC, but contain a relevant proportion of the posterior of some later observed dataset, thus making the results of prior SBC too optimistic for that dataset. Second, the time required for running the inference on each SBC iteration is often non-trivial and one quickly needs to consider the tradeoff between their computational budged and the power of the calibration checking. Third, as visualized in Figure 1, even with unlimited computational resources the result of prior SBC does not guarantee calibration within a specific subregion of the inspected parameter space. There can always be another region with an equally strong but reverse effect on the overall distribution of the PIT values. As proposed by Modrák et al., (2023), one solution to this last problem is to employ additional test quantities using functions of parameters and non-monotonic transformations (such as folded rank statistics), that might detect these problematic regions, but coming up with these test quantities can be difficult.
@@
2 Posterior simulation-based calibration checking – SBC conditioned on observed data
Figure 2: Three iterations of prior SBC for a model of a normal distribution with a known standard deviation. Estimate of the mean parameter before (dashed line) and after (solid lines) conditioning on predictive draws. The PIT value of each prior draw (blue dot) equals the respective shaded blue area ($0.03$, $0.43$, and $0.97$ respectively).
@@ 2 Posterior simulation-based calibration checking – SBC conditioned on observed data
Weakly informative priors are commonly used in Bayesian inference due to their desirable properties in real-life applications (Gelman et al.,, 2017). However, this class of priors can cause serious issues for interpreting the results of prior SBC. Even though a large portion of the prior mass may lie on a region of the parameter space where the inference works well, smaller regions may exist where the inference can fail. On the one hand, if large enough, these issues may be detected through prior SBC. The modeller could then try to modify the priors or the model to reduce the prior probability of the problematic parameter values. On the other hand, if one already has access to the data that the model will later be used to run inference on, the assessment of the inference algorithm could be focused on the region of the parameter space around the posterior. For inference on the given data, this is arguably the most important region to ensure proper calibration for.
@@
Using the sequential Bayesian updating rule, one can view the posterior as the new prior and perform the entire SBC procedure conditional on the observed data $y_{obs}$. In other words, by conditioning the joint distribution in 2 on the observed data,
@@
we see that a sample from the old posterior $\theta^{\prime}_{i}\sim\pi(\theta\,|\,y_{obs})$, and $\theta^{\prime\prime}\sim\pi(\theta\,|\,y_{i},y_{obs})$ drawn from an augmented posterior, need to have the same distribution given a prediction $y_{i}\sim\pi(y\,|\,\theta^{\prime}_{i})$. Thus, we operationalize the approach for data conditioned SBC by drawing $\theta^{\prime}_{i}\sim\pi(\theta\,|\,y_{\text{obs}})$, generating posterior predictions $y_{i}\sim\pi(y\,|\,\theta^{\prime}_{i})$, and then using the inference algorithm to be validated to draw from the new posterior $\theta^{\prime\prime}_{1},\dots\theta^{\prime\prime}_{S}\sim\pi(\theta\,|\,y_{i},y_{\text{obs}})$.
@@
The validity of the algorithm can be assessed by testing the uniformity of PIT values, $u_{i}=p(\theta^{\prime\prime}_{i}<\theta^{\prime}_{i}\,|\,y_{i},y_{obs})$, this time computed for the original posterior draws, $\theta^{\prime}_{i}$, with regard to the (thinned) augmented posterior draws, $\theta^{\prime\prime}_{i,1},\dots,\theta^{\prime\prime}_{i,N}$. We call this variant of SBC using posterior predictive draws posterior SBC.
@@
When using prior SBC for validating the inference algorithm, drawing parameter values generatively from the prior is usually assumed to be trivial. In posterior SBC, we need to sample from two different posteriors, which would usually achieved using the same algorithm that is being evaluated. Although we don’t have exact draws from one of the posteriors, the consistency of Equation 3 does not hold if the evaluated algorithm is not successfully sampling from both posteriors, leading to non-uniform PIT values in posterior SBC.

### v2

@@ 1 Introduction
(a) Two cancelling biases
(b) One problematic region
Figure 1: Two conceptual illustrations of a model parameter space, with gray area denoting the prior. In (a), the colored regions are potential posteriors with bias in opposite directions, while the inference is well calibrated for parameter values outside these regions. Prior SBC will not show calibration issues due to cancellation of biases, while posterior SBC would indicate issues for posteriors intersecting the colored regions. In (b), only one region is problematic. Now prior SBC would indicate calibration issues, while posterior SBC would show that the inference is calibrated outside the colored region. In both cases posterior SBC has the benefit of focusing on the region, which matters given the observed dataset.
@@
To ensure trust on the results of the inference, different forms of calibration checking play important roles in the Bayesian workflow for model building (Gelman et al.,, 2020). Simulation-based calibration checking (SBC; Cook et al.,, 2006; Talts et al.,, 2020; Modrák et al.,, 2023) is one of these methods and validates the chosen posterior inference algorithm as well as checks for inconsistencies between the model implementation and a possibly separate implementation of the data generating process. Probabilistic programming software (Štrumbelj et al.,, 2024) provide probabilistic programming languages to make it easier to implement probabilistic models, and inference engines that implement various inference algorithms. As the probabilistic programming software and algorithms for probabilistic modelling advance, the threshold for specifying increasingly complicated models gets lowered. However, as the complexity of the models increases, so does the probability of human error when writing the code, which implements the model. In addition to such model code implementation mistakes, the inference algorithms used for obtaining posterior approximations—even when correctly coded—can work well for one posterior, but suffer from computational difficulties for another. By simply changing the observed data, one may obtain vastly different posterior geometries, favoring one implementation or algorithm over another.
SBC aims to check that the inference is calibrated, but miscalibration can be caused either a mistake in the model implementation code, a mistake in inference algorithm implementation code, or significant bias in the inference algorithm. The algorithm is considered to be calibrated, if the probability integral transformation (PIT) of the parameters is uniformly distributed, when transformed with respect to the posterior conditional on data generated using those parameters. We review SBC and the definition of calibration in more detail in Section 2.
In this paper we assume that the both the code for implementing the model, and the inference algorithm implementation are correct and focus on examples were the inference algorithm may produce significantly biased results with finite computation time. Even when the inference algorithm is implemented correctly, it may produce significantly biased inference for some posteriors. For example, many Markov chain Monte Carlo (MCMC) and variational inference approaches have challenges with funnel shaped posteriors (see, e.g. Neal,, 2003; Yao et al.,, 2018), which are common in case of hierarchical models. The original SBC is limited to checking for calibration on the whole joint distribution of parameters and observations under the specified priors, and thus we call it prior SBC. Specifying prior distributions that accurately present the available prior information is often difficult. In some cases, the prior may have significant probability mass in parts of the parameter space that generate such data for which the inference algorithm produce significantly biased inference (in finite time). For example, the data generated from a hierarchical model with certain type of parameter values can lead to a funnel shaped posterior which has a challenging geometry for many inference algorithms to explore. Figure 1 illustrates two cases: a) in which the prior space includes two subspaces with canceling biases and SBC indicates that the calibration over the prior is good, and b) in which the prior space includes one subspace with biased inference causing SBC to indicate that the calibration over the prior is not good. However, after observing some data we are not interested in calibration over the prior, but calibration within the region of the posterior.
In this paper, we introduce posterior SBC, a variant of SBC using the self-consistency of Bayesian posterior distribution to validate the model implementation and inference algorithm when conditioned on some fixed set of data. Figure 1 illustrates two cases where, if the posterior is much more concentrated than the prior, posterior SBC checks the calibration of the inference for those regions of the parameter space where it matters and does not care how well the inference algorithm works elsewhere. Prior SBC is useful for algorithm and software developers, while posterior SBC is useful for modellers. As the number of modellers is several orders of magnitude bigger than the number of developers, we expect the posterior SBC to become more popular than prior SBC.
@@
We next review the method of simulation-based calibration as it has traditionally been operationalized. In Section 3, we introduce posterior SBC. In Section 4 we illustrate the differences between prior and posterior SBC through three case studies including a simple hierarchical model, Lotka-Volterra model, and a drift-diffusion model. In the first two examples we use MCMC and in the last example we use amortized Bayesian inference for which there are no similar convergence diagnostics as for MCMC. Finally, in Section 5 we summarize the contributions of the paper and provide an outlook for future research.
2 Simulation-based calibration checking
Let us consider a Bayesian model of the joint distribution of the data $y$ and parameters $\theta$,
$\pi(y,\theta)=\pi(y\,|\,\theta)\pi(\theta),$
where $\pi(\theta)$ is the prior distribution of the parameters and $\pi(y\,|\,\theta)$ the likelihood of the data.
Cook et al., (2006) propose a simulation-based calibration checking method for validating Bayesian inference software designed to fit the model. The key to this method is the following self-consistency property of Bayesian posterior distributions. Let $\theta^{\prime}\sim\pi(\theta)$, and $y^{\prime}\sim\pi(y\,|\,\theta^{\prime})$. Now the pair $(y^{\prime},\theta^{\prime})$ represents a draw from the joint distribution $\pi(y,\theta)$, and therefore $\theta^{\prime}$ is a draw from the posterior distribution, $\pi(\theta\,|\,y^{\prime})$. This self-consistency property, linking the prior, prior predictive, and posterior distributions, can be summarized with the following SBC equality:
$\pi(y^{\prime},\theta^{\prime},\theta^{\prime\prime})=\pi(\theta^{\prime\prime}\,|\,y^{\prime})\pi(y^{\prime}\,|\,\theta^{\prime})\pi(\theta^{\prime})=\pi(\theta^{\prime\prime}\,|\,y^{\prime})\pi(\theta^{\prime}\,|\,y^{\prime})\pi(y^{\prime}),$ (1)
where $\theta^{\prime\prime}\sim\pi(\theta\,|\,y^{\prime})$ is a posterior draw conditioned on $y^{\prime}$.
Given an observation $y^{\prime}$ and the model implementation, the inference algorithm should be able to sample from the posterior, $\theta^{\prime\prime}\sim\pi(\theta\,|\,y^{\prime})$. Therefore, to verify that the inference algorithm is able to accurately sample from the posterior, we should assess whether the posterior draw $\theta^{\prime\prime}$ and the prior draw $\theta^{\prime}$, conditional on $y^{\prime}\sim\pi(y\,|\,\theta^{\prime})$, are from the same distribution.
@@ 2 Simulation-based calibration checking
To assess whether $\theta^{\prime}$ and $\theta^{\prime\prime}$ are from the same distribution conditional on $y$, Cook et al., (2006) repeatedly draw parameter vectors $\theta^{\prime}_{i}$ from the prior $\pi(\theta)$, generate data from the observation model $y_{i}\sim\pi(y_{i}\,|\,\theta^{\prime}_{i})$, and then use the algorithm to be validated to sample $\theta^{\prime\prime}_{i,1},\ldots,\theta^{\prime\prime}_{i,S}\sim\pi(\theta^{\prime\prime}\,|\,y_{i})$. The authors then propose to compute empirical probability integral transform (PIT) values,
$u_{i}=p(\theta_{i}^{\prime\prime}<\theta_{i}^{\prime}\,|\,y_{i}),$ (2)
which prove to be uniformly distributed as $S\to\infty$. The process is repeated for $i=1,\ldots,N$ and the empirical PIT values are used for testing. Figure 2 shows three iterations of prior draws and their respective PIT relative to the corresponding posterior distributions. Cook et al., (2006) further propose the use of the $\chi^{2}$-test for the inverse of the normal cumulative distribution function (CDF) of the empirical PIT values. The inference is said to be calibrated if the PIT values pass a uniformity test.
Figure 2: An illustration of how PIT values are computed in prior SBC. The model is a normal distribution $\mathcal{N}(\theta,\sigma)$ with a known standard deviation $\sigma$. The dashed line shows the prior distribution $p(\theta)$, and solid lines show three posteriors $p(\theta|y_{i})$ conditioning on the predictive draws $y_{i}\sim\pi(y_{i}\,|\,\theta^{\prime}_{i})$. The PIT value of each prior draw $\theta^{\prime}_{i}$, shown as blue dots, equals the respective shaded blue area ($0.03$, $0.43$, and $0.97$ respectively). In practice the PIT values (2) are computed using Monte Carlo draws from the posteriors $\theta^{\prime\prime}_{i,1},\ldots,\theta^{\prime\prime}_{i,S}\sim\pi(\theta^{\prime\prime}\,|\,y_{i})$.
@@
Instead of only assessing the uniformity of the ranks of individual parameters, Modrák et al., (2023, Sections 3.3–3.4) propose the use of the joint log-likelihood $\log p(y\,|\,\theta)$ as a test statistic for assessing the calibration. That is, $p(y\,|\,\theta^{\prime})$ and $p(y\,|\,\theta^{\prime\prime})$ are used instead of $\theta^{\prime}$ and $\theta^{\prime\prime}$ when computing PIT values with (2). This has the benefit of being a joint function of both data and all parameters. Using the data in the test quantity allows detecting calibration issues rising from the model partially ignoring the data. Modrák et al., (2023, Sections 3.4 and 4.3) demonstrate how test quantities based on only parameters can not detect if the model completely ignores all data (the posterior is the prior) or part of the data (e.g. the first data point), but using the log-likelihood as a test quantity indicates the problem. Another advantage of using the log-likelihood as a test quantity is in checking calibration of models with high-dimensional parameter spaces, where modelers might not be particularly interested in gauging the calibration of all the individual parameters, and using the log-likelihood as a test quantity is a natural choice for a joint function of model parameters. Using the joint log-likelihood also avoids the need of correcting for multiple comparisons from simultaneously assessing the calibration of many individual parameters.
Performing SBC on the full prior parameter space has three fundamental drawbacks. First, as shown in the case study of Section 4.1, it is possible that only some parameter values generate such data that lead to a posterior from which the inference method is not able to sample faithfully. If such parameter values occur in a small region of the parameter space, and we average over a much wider prior, the effect in the calibration checking can be minor and prior SBC may miss the miscalibration completely. If such parameter values have high probability mass under the posterior, then posterior SBC shows the miscalibration.
@@
Second, the time required for running the inference on each SBC iteration is often non-trivial and one needs to consider the tradeoff between their computational budget and the power of the calibration checking. Third, as visualized in Figure 1, even with unlimited computational resources the result of prior SBC does not guarantee calibration within a specific subregion of the inspected parameter space. There can always be another region with an equally strong but reverse effect on the overall distribution of the PIT values. As proposed by Modrák et al., (2023), one solution to this last problem is to employ additional test quantities using functions of parameters and non-monotonic transformations (such as folded rank statistics), that might detect these problematic regions, but coming up with these test quantities can be difficult.
@@
3 Posterior simulation-based calibration checking – conditioned on observed data
@@
Weakly informative priors are commonly used in Bayesian inference due to their desirable properties in real-life applications (Gelman et al.,, 2017). However, this class of priors can cause serious issues for interpreting the results of prior SBC. Even though a large portion of the prior mass may lie on a region of the parameter space where the inference works well, smaller regions may exist where the inference method can fail to faithfully present the posterior. On the one hand, if large enough, these issues may be detected through prior SBC. The modeller could then try to modify the priors or the model to reduce the prior probability of the problematic parameter values. On the other hand, if one already has access to the data that the model will later be used to run inference on, the assessment of the inference algorithm could be focused on the region of the parameter space around the posterior. For inference on the given data, this is arguably the most important region to ensure proper calibration for.
@@
Using the sequential Bayesian updating rule, one can view the posterior as the new prior and perform the entire SBC procedure conditional on the observed data $y_{obs}$. In other words, by expanding the joint distribution in (1) to include both the observed data and the predictive draw,
@@
we see that a sample from the old posterior, $\theta^{\prime}_{i}\sim\pi(\theta\,|\,y_{obs})$, and $\theta^{\prime\prime}\sim\pi(\theta\,|\,y_{i},y_{obs})$ drawn from an augmented posterior, need to have the same distribution given the prediction $y_{i}\sim\pi(y\,|\,\theta^{\prime}_{i})$. Thus, we operationalize the approach for data conditioned SBC by drawing $\theta^{\prime}_{i}\sim\pi(\theta\,|\,y_{\text{obs}})$, generating posterior predictions $y_{i}\sim\pi(y\,|\,\theta^{\prime}_{i})$, and then using the inference algorithm to be validated to draw from the new posterior $\theta^{\prime\prime}_{1},\dots\theta^{\prime\prime}_{S}\sim\pi(\theta\,|\,y_{i},y_{\text{obs}})$.
@@
The calibration of inference can be checked by testing the uniformity of the PIT values, $u_{i}=p(\theta^{\prime\prime}_{i}<\theta^{\prime}_{i}\,|\,y_{i},y_{obs})$, this time computed for the original posterior draws, $\theta^{\prime}_{i}$, with regard to the (thinned) augmented posterior draws, $\theta^{\prime\prime}_{i,1},\dots,\theta^{\prime\prime}_{i,N}$. We call this variant of SBC using posterior draws posterior SBC.
@@
When using prior SBC for validating the inference algorithm, drawing parameter values generatively from the prior is usually assumed to be trivial. In posterior SBC, we need to sample from two different posteriors, which would usually be achieved using the same algorithm that is being evaluated. Posterior SBC does not require that we are able to get draws from the true posterior. If the sampling fails for either the original posterior or for the augmented posteriors, the SBC equality in Equation 3 does not hold, leading to non-uniform PIT values in posterior SBC. While the Bayesian updating is consistent when more data is observed, it is very unlikely that we would encounter such biased inference that would be consistent when conditioned on more data.
The posterior SBC looks similar to some other methods using draws from the posterior predictive distribution. Posterior predictive checking (PPC; Box,, 1980; Rubin,, 1984; Gelman et al.,, 1996, 2013; Gabry et al.,, 2019) compares draws from the posterior predictive distribution to the original data. PPC uses the data twice which can lead to optimistic checking unless the test statistic is completely ancillary which is unlikely especially for more flexible models. Parameter recovery experiments generate data from the model using some feasible parameter values, and the posterior is compared to those parameter values visually without a formal test. The crucial part of posterior SBC is to use the chain rule and compare the draws from the original posterior to the augmented posterior which is conditioned on both the original data and the posterior predictive draws.
A reader may wonder why do we need posterior SBC, when sampling from the original posterior with Markov chain Monte Carlo can be diagnosed with convergence diagnostics such as $\widehat{R}$ (Vehtari et al., 2021a, ) and divergences (Betancourt,, 2017). Vehtari et al., 2021a (, Appendix C) demonstrate how $\widehat{R}$ fails to diagnose bias in the case of a funnel shaped posterior and very long chains. While in that example Hamiltonian Monte Carlo specific divergence diagnostic did work, it has also been observed to miss convergence issues. Furthermore, additional diagnostics similar to $\widehat{R}$ and the divergence diagnostic are not available for all MCMC methods. What is more, while convergence diagnostics can indicate issues, they do not indicate the direction or magnitude of the induced bias. For example, it is known that there are also false positive divergence warnings, and in such cases posterior SBC can be used to measure the amount of potential bias. While the posterior SBC is operationalized by sampling from the posterior, the inference algorithms tested do not need to be (Markov chain) Monte Carlo methods. Any inference algorithm can be tested as long as we can also get draws from the approximated posterior. (Yao et al.,, 2018) demonstrate the challenges of diagnosing the reliability of variational inference in high dimensions and use prior SBC to check variational inference by drawing from the variational approximation. Schad et al., (2023) discuss problems in estimating Bayes factors and use prior SBC for checking computation of Bayes factors with bridge sampling and Savage-Dickey method. Considering (Yao et al.,, 2018) and Schad et al., (2023) did consider reliability of computation given the observed data, it would have been sensible to use posterior SBC, instead. In Section 4.3 we illustrate the use posterior SBC for amortized Bayesian inference, for which before this there were no practical inference checking tools.